# Targeted Error Correction in Knowledge Distillation: Small Language Models Surpass GPT

### v1

@@ Abstract
We introduce an Analyze–Revise–Finetune (ARF) pipeline that enables smaller open-source language models (LLMs) to surpass substantially larger proprietary models in customer service summarization tasks. The pipeline first analyzes and categorizes common errors in summaries produced by a teacher model (GPT-3.5), then performs a targeted revision using a compact editor model (Llama 3.1 70B) to generate high-quality, refined training data. Fine-tuning a smaller student model (Llama 3.1 8B) on this refined data resulted in superior summarization performance compared to GPT-3.5. The ARF pipeline improves cost efficiency and data privacy while maintaining competitive accuracy, illustrating a generalizable framework for enhancing open-source LLMs across diverse downstream applications.
Keywords Automatic error correction $\cdot$ LLM fine-tuning $\cdot$ Summarization
@@ 1 Introduction
In this paper, we introduce an "Analyze-Revise-Finetune" (ARF) pipeline (illustrated in Figure 1) designed to enhance the performance of smaller, open-source LLMs (student models) so they surpass the capabilities of larger, proprietary teacher models. ARF comprises three primary steps: 1) systematic identification of prevalent error patterns in auto-generated texts, 2) targeted error correction using a separate editor LLM, and 3) leveraging the corrected texts to fine-tune a smaller open-source student LLM. Unlike previous studies that either exclusively depend on teacher-generated pseudo-labels or iterative self-refinement, ARF incorporates a targeted error correction step based on an error-focused evaluation scheme that facilitates a systematic analysis of teacher-generated texts to pinpoint prevalent errors in the training data. By specifically targeting these common errors for correction, we remove the need for repetitive critic-and-revision cycles. Furthermore, targeted error correction is inherently simpler than full critic-revision, allowing the editor LLM to be smaller and more cost-effective than the teacher model. By fine-tuning the student LLM using the revised training data, the student model learns effectively from the teacher without inheriting its frequent errors, unlike methods that uncritically adopt all teacher-generated pseudo-labels.
We demonstrate the effectiveness of ARF pipeline in a customer service interaction summarization task currently served by GPT-3.5. Initially, we define comprehensive evaluation criteria to systematically identify and classify specific error patterns present in GPT-3.5-generated summaries. Utilizing these criteria, we identify common errors prevalent across summaries. Subsequently, we employ a smaller editor model (Llama 3.1 70B) to specifically revise these identified errors, thus creating a refined training dataset. Finally, we fine-tune a smaller open-source model (Llama 3.1 8B) on this revised dataset, demonstrating that our fine-tuned model successfully replicates GPT-3.5’s functionality while significantly reducing frequent errors, ultimately outperforming GPT-3.5 in summary generation.
@@ 2 Related Work
Many studies leverage larger models ("teacher" models) to generate pseudo-labels for training smaller models ("student" models). Hinton et al. [1] introduced a knowledge distillation method where the teacher model’s output probabilities serve as soft targets for the student model training. This approach enables the student to capture nuanced predictions from the teacher, effectively consolidating knowledge from an ensemble of models into a single model. Gu et al. [2] extended knowledge distillation to large language models (LLMs), employing reverse Kullback-Leibler divergence (KLD) instead of the conventional forward KLD, comparing the student LLM’s output distribution directly with that of the teacher. Sanh et al. [3] developed DistilBERT, which achieves 97% of BERT’s language understanding capability while being 40% smaller by minimizing the KL divergence between student and teacher softmax outputs. Unlike these approaches, which typically treat all teacher-generated outputs equally without any error analysis or correction, ARF provides a focused approach where common errors in teacher-generated texts are explicitly identified and corrected, thereby ensuring higher-quality pseudo-labels.
Recent research has investigated iterative refinement techniques involving LLMs, where a model evaluates and improves its own generated outputs. "Self-Refine"[4] is a framework that allows an LLM to iteratively critique and refine its responses autonomously without external supervision or additional training data. Similarly, LLM-AUGMENTER[5] addresses hallucinations in LLM outputs by integrating external knowledge and iterative prompt revision based on feedback from utility functions. Bai et al. [6] proposed Constitutional AI, a self-supervised framework that utilizes guiding principles (a "constitution") to critique and iteratively revise an LLM’s outputs, ultimately using these revised outputs to train a safer model. In contrast to these iterative self-refinement methods, ARF emphasizes targeted correction of the most common errors identified through systematic analysis. Additionally, while previous methods generally rely on a single LLM for both output generation and self-correction, ARF separates initial generation (performed by teacher LLM) from subsequent error correction (performed by editor LLM). This separation enables the use of a smaller and more cost-efficient editor model, achieving high-quality revisions with lower inference costs.
@@
In 2024, eBay deployed an LLM-based summarization system aimed at enhancing the productivity of customer service agents (referred to as teammates). To effectively assist customers, teammates must first understand the customer’s issue or request by reviewing any previous interactions the customer had with eBay. The primary role of the summarization system is to succinctly summarize these prior interactions, highlighting essential information while removing any superfluous details, enabling teammates to efficiently and promptly take appropriate next steps.
This summarization system processes textual input from customer interactions originating from four distinct communication channels: bot-customer chat, teammate-customer chat, emails, and webforms. Before summarization, the system automatically redacts personal information from the input content. The sanitized text is then sent to GPT-3.5, which generates concise summaries presented as bullet points within the Service Desktop interface. To facilitate continuous improvement, a feedback mechanism allowing teammates to give a thumbs-up or thumbs-down to indicate summary quality is integrated into the Service Desktop. Since its initial launch in February 2024, the summarization system has maintained a thumbs-down feedback rate below 5%.
We fine-tune a compact open-source LLM for customer service summarization with the goal of replacing GPT-3.5 in production. This approach substantially reduces inference costs and mitigates privacy risks inherent to proprietary models. Our experiments focus on two communication channels—bot–customer chats (BotChat) and webforms (WebForm)—which account for most high-volume, text-based interactions in the production environment.
@@ 4 Error-Focused Summary Evaluation Scheme
To identify opportunities for further enhancing the quality of the production summarizer and to support AI governance efforts, we retrieved a representative sample of input content and corresponding summaries from production data. These samples were then manually reviewed by a team of subject matter experts and the AI governance team. We collaborated closely with this team to develop a structured summary evaluation scheme aimed at ensuring consistency in the review process and facilitating systematic identification of recurring issues and potential areas for improvement. This evaluation scheme is centered around the identification of specific "errors" within summaries, categorized based on predefined error types derived from an initial data analysis.
Errors are grouped into seven primary categories: Content, Entities, Data Elements, Customer Type, Unnecessary Information, Inferred Sentiment, and Language. Each category is further divided into detailed sub-labels specifying the precise nature of errors. The Content category covers fundamental summary issues such as missing critical details, inaccuracies, or fabricated content. The Entities category addresses problems related to missing or inaccurately represented key identifiers, including item numbers, names, order numbers, and other transaction-related details. Data Elements encompasses inaccuracies in structured information, such as dates, that are not captured by the Entities category. Customer Type focuses on incorrectly identifying or omitting the customer’s role (buyer or seller), a specific error relevant only to WebForm data. Unnecessary Information flags the inclusion of irrelevant or redundant details, including unnecessary bot interactions, courteous engagements, or repeated information. Inferred Sentiment targets unwarranted emotional inferences about the customer that are unsupported by the original interaction. Lastly, Language captures errors associated with incorrect or missing language identification, as well as translation inaccuracies. This comprehensive categorization framework supports a consistent, precise, and detailed evaluation of summary quality. A complete taxonomy of summary error types is available in Table 5 in the appendix.
Summaries are rated on a 1-to-5 scale according to the number and severity of errors identified. A rating of 1 indicates numerous or critical errors significantly impacting the overall understanding of the situation and the subsequent customer service. A rating of 2 denotes summaries containing one or two errors. A summary receives a rating of 3 if it contains only a minor error that does not affect the overall understanding. Summaries rated 4 are considered acceptable, with no distinct errors but potentially lacking perfect conciseness. Finally, a rating of 5 represents summaries that are flawless, containing no errors or issues.
We utilize this evaluation scheme to assess the quality of summaries generated by various LLMs and to compare the performance of different fine-tuned models (Section 6). Additionally, we use the predefined error types in the scheme to systematically analyze GPT-3.5-generated summaries and identify the most common errors (Section 5.2).
@@ 5.1 Data
Due to the presence of personally identifiable information (PII) in customer interaction logs, we developed a secure data retrieval pipeline that downloads production data samples from the RNPCI zone and anonymizes all PII by substituting them with artificial fake values. Using this pipeline, we collected 10,000 samples per communication channel to serve as training data. Specifically for WebForm data, additional preprocessing steps were applied to parse the XML structure and extract only the information pertinent to customer service interactions. Furthermore, we retrieved separate test and development sets comprising 200 samples each per channel. Lastly, smaller datasets consisting of 100 BotChat and 68 WebForm samples were obtained to facilitate detailed analysis of prevalent error types in production-generated summaries.
@@ 5.2 Systematic error analysis
To identify the most frequent error types occurring in summaries produced by GPT-3.5, a team of subject matter experts reviewed production summaries—specifically 100 BotChat samples and 68 WebForm samples. Each summary was evaluated for errors based on the predefined error types outlined in our summary evaluation scheme (Table 5). After completing the evaluation, we determined the most common error types present in the summaries from each channel. Table 1 summarizes these key error types.
Table 1: Major error types across different channels
Bot Chat WebForm
Error Type Count Error Type Count
unn_content_requests_agent 30 (38.0%) unn_content_redundant 14 (29.2%)
sentiment_inferred_frustrated 19 (24.0%) customer_type_inaccurate 7 (14.6%)
content_missing 13 (16.5%) unn_content_webform_email_copy 4 (8.3%)
nothing_to_summarize 6 (7.6%) entity_missing_item_number 3 (6.3%)
sentiment_inferred_not_frustrated 5 (6.3%) content_inaccurate 3 (6.3%)
all errors 79 (100%) all errors 48 (100%)
@@ 5.3 Automatic error-correction
From the error types identified through systematic analysis (Section 5.2), we selected four frequent and manageable errors for automated correction. For BotChat, we targeted the errors unn_content_requests_agentunn\_content\_requests\_agent, referring to unnecessary mentions in summaries of customers requesting human assistance, and sentiment_inferred_frustratedsentiment\_inferred\_frustrated, referring to inferred customer frustration not explicitly stated by the customer. For WebForm summaries, we selected the errors unn_content_redundantunn\_content\_redundant, denoting unnecessary redundant information, and unn_content_webform_email_copyunn\_content\_webform\_email\_copy, indicating the unnecessary inclusion in summaries of customer requests for email copies of their submitted webforms.
We employed Llama 3.1 70B as the editor LLM for performing these error corrections. Being an open-source model, Llama 3.1 70B provides reduced privacy risks. Correction prompts designed for each error type are depicted in Figure 2. Specifically, the prompt addressing sentiment_inferred_frustratedsentiment\_inferred\_frustrated was designed broadly to remove all references to sentiment rather than just inferred frustration. This approach does not negatively impact summary quality since sentiment information is considered non-essential according to our evaluation criteria (Section 4).
We processed all 20,000 training samples across both channels using a cascading sequence of error-correction prompts. For BotChat summaries, corrections were first applied to remove sentiment_inferred_frustratedsentiment\_inferred\_frustrated, followed by unn_content_requests_agentunn\_content\_requests\_agent. For WebForm summaries, corrections began with unn_content_webform_email_copyunn\_content\_webform\_email\_copy. Subsequently, the prompt addressing unn_content_redundantunn\_content\_redundant was applied to both channels, under the assumption that increased conciseness would benefit both datasets. We denote the dataset prior to applying the unn_content_redundantunn\_content\_redundant fix as revision version 1 (r1), and the dataset after applying this correction as revision version 2 (r2).
Prompt to remove unn_content_requests_agentunn\_content\_requests\_agent error from BotChat summaries Read the following summary of a chat transcript between an eBay customer and an automatic service agent. Rewrite the summary to remove any information about the customer wanting or requesting to connect with a human agent. If there’s no other content in the summary other than customer wanting or requesting to connect with a human agent, say "<ul><li>nothing to summarize<li></ul>". Keep the original format of the summary that is unordered list items in HTML. Do not add anything else other that the revised summary.
Prompt to remove sentiment_inferred_frustratedsentiment\_inferred\_frustrated error from BotChat summaries Read the following summary of a chat transcript between an eBay customer and an automatic service agent. Rewrite the summary to remove any information about the customer’s emotion such as being frustrated, any information about the reason of the emotion, and any information about how the customer expressed their emotion. If there’s no other content in the summary other than customer’s emition, say "<ul><li>nothing to summarize<li></ul>". Keep the original format of the summary that is unordered list items in HTML. Do not add anything else other that the revised summary.
Prompt to remove unn_content_webform_email_copyunn\_content\_webform\_email\_copy error from WebForm summaries Read the following summary of a WebForm submitted by an eBay customer. Rewrite the summary to remove any information about the customer wanting or requesting a copy of the communication to their email. Do not add anything else other that the revised summary.
Prompt to remove unn_content_redundantunn\_content\_redundant for both BotChat and WebForm summaries Read the following summary. Rewrite the summary to remove any information that is repeatedly stated. Do not remove any important information such as Item Ids, Return Ids, Case Ids, etc or customer’s role as a buyer or a seller. Keep the original format of the summary that is unordered list items in HTML. Do not add anything else other that the revised summary.
Figure 2: Prompts for summary error correction
@@ 5.4 LLM fine-tuning
We evaluated multiple Llama and Lilium models of different sizes as student LLMs. Each model was fine-tuned independently using three dataset versions: the original dataset (org), revision version 1 (r1), and revision version 2 (r2). Instruction-tuning samples were generated using the same summary generation prompts currently utilized by the production GPT-3.5 model. Fine-tuning was performed with Low-Rank Adaptation (LoRA) [7], applying parameters $alpha=16$, $dropout=0.1$, and $r=8$ across all linear modules.
Fine-tuning of the Lilium 2 mix 7B and Lilium 2 13B models was conducted using Athena, while other models utilized between 4 and 8 A100 GPUs depending on model size. During inference, we used the same prompts as those employed in the production environment for all experiments.
@@ 6 Evaluation
Specifically, we used GPT-4 Turbo as the judge LLM and performed prompt tuning separately for each communication channel to output ratings on a 1-to-5 scale, based on the original content and corresponding generated summaries. Our automatic evaluation system demonstrated strong correlation with human ratings when benchmarked against other advanced LLM-as-judge methodologies [8, 9]. For the BotChat channel, the correlation was measured with a Spearman’s $\rho$[10] of 0.6663 and Kendall’s $\tau$[11] of 0.6005. For the WebForm channel, Spearman’s $\rho$ was 0.5674, and Kendall’s $\tau$ was 0.6364.
Our primary evaluation metric is the mean score of automatic ratings provided by our evaluators on the test dataset.

### v2

# Error-Aware Knowledge Distillation via Targeted Revision for Customer-Service Summarization

@@ Abstract
We introduce an Analyze–Revise–Finetune (ARF) pipeline that enables smaller open-source language models (LLMs) to surpass substantially larger proprietary models in customer service summarization tasks. The pipeline first analyzes and categorizes common errors in summaries produced by a teacher model (GPT-3.5), then performs a targeted revision using a compact editor model (Llama 3.1 70B) to generate high-quality, refined training data. Fine-tuning smaller student models (e.g., Llama 3.1 8B, QWen3 4B) on this refined data resulted in superior summarization performance compared to GPT-3.5. The ARF pipeline improves cost efficiency and data privacy while maintaining competitive accuracy, illustrating a generalizable framework for enhancing open-source LLMs across diverse downstream applications.
Keywords LLM fine-tuning $\cdot$ Knowledge distillation $\cdot$ Summarization
@@ 1 Introduction
In this paper, we introduce an Analyze-Revise-Finetune (ARF) pipeline (illustrated in Figure 1) designed to enhance the performance of smaller, open-source LLMs (student models) so they surpass the capabilities of larger, proprietary teacher models. ARF comprises three primary steps: 1) systematic identification of prevalent error patterns in auto-generated texts, 2) targeted error correction using a separate editor LLM, and 3) leveraging the corrected texts to fine-tune a smaller open-source student LLM. Unlike previous studies that either exclusively depend on teacher-generated pseudo-labels or iterative self-refinement, ARF incorporates a targeted error correction step based on an error-focused evaluation scheme that facilitates a systematic analysis of teacher-generated texts to pinpoint prevalent errors in the training data. By specifically targeting these common errors for correction, we remove the need for repetitive critic-and-revision cycles. Furthermore, targeted error correction is inherently simpler than full critic-revision, allowing the editor LLM to be smaller and more cost-effective than the teacher model. By fine-tuning the student LLM using the revised training data, the student model learns effectively from the teacher without inheriting its frequent errors, unlike methods that uncritically adopt all teacher-generated pseudo-labels.
We demonstrate the effectiveness of ARF pipeline in a real-world customer service interaction summarization task that was served by GPT-3.5. Initially, we define comprehensive evaluation criteria to systematically identify and classify specific error patterns present in GPT-3.5-generated summaries. Utilizing these criteria, we identify common errors prevalent across summaries. Subsequently, we employ an open source editor model (Llama 3.1 70B) to specifically revise these identified errors, thus creating a refined training dataset. Finally, we fine-tune smaller open-source models on this revised dataset, demonstrating that our fine-tuned models successfully replicate GPT-3.5’s functionality while significantly reducing frequent errors, ultimately outperforming GPT-3.5 in summary generation.
@@ 2 Related Work
Many studies leverage larger models (teacher models) to generate pseudo-labels for training smaller models (student models). Hinton et al. [5] introduced a knowledge distillation method where the teacher model’s output probabilities serve as soft targets for the student model training. This approach enables the student to capture nuanced predictions from the teacher, effectively consolidating knowledge from an ensemble of models into a single model. Gu et al. [4] extended knowledge distillation to large language models (LLMs), employing reverse Kullback-Leibler divergence (KLD) instead of the conventional forward KLD, comparing the student LLM’s output distribution directly with that of the teacher. Sanh et al. [14] developed DistilBERT, which achieves 97% of BERT’s language understanding capability while being 40% smaller by minimizing the KL divergence between student and teacher softmax outputs. Unlike these approaches, which typically treat all teacher-generated outputs equally without any error analysis or correction, ARF provides a focused approach where common errors in teacher-generated texts are explicitly identified and corrected, thereby ensuring higher-quality pseudo-labels.
Recent research has investigated iterative refinement techniques involving LLMs, where a model evaluates and improves its own generated outputs. Self-Refine [10] is a framework that allows an LLM to iteratively critique and refine its responses autonomously without external supervision or additional training data. Similarly, LLM-AUGMENTER [12] addresses hallucinations in LLM outputs by integrating external knowledge and iterative prompt revision based on feedback from utility functions. Bai et al. [1] proposed Constitutional AI, a self-supervised framework that utilizes guiding principles (a "constitution") to critique and iteratively revise an LLM’s outputs, ultimately using these revised outputs to train a safer model. In contrast to these iterative self-refinement methods, ARF emphasizes targeted correction of the most common errors identified through systematic analysis. Additionally, while previous methods generally rely on a single LLM for both output generation and self-correction, ARF separates initial generation (performed by teacher LLM) from subsequent error correction (performed by editor LLM). This separation enables the use of a smaller and more cost-efficient editor model, achieving high-quality revisions with lower inference costs.
@@
We base our study on a real-world use case using data from a large e-commerce platform. The platform launched an LLM-based summarization system to improve customer service agent productivity. In typical support workflows, agents must understand a customer’s issue by reviewing prior interactions across multiple communication channels. These interactions can span multiple sessions and contain lengthy, unstructured dialogue, making it time-consuming to reconstruct context before resolving the issue.
The system summarizes customer interactions from four channels: bot-customer chat (BotChat), teammate-customer chat (TeammateChat), emails (Email), and webforms (Webform). To enable continuous improvement, it includes a feedback mechanism where teammates provide thumbs-up or thumbs-down ratings. Since its launch in February 2024, the thumbs-down rate has remained below 5
We fine-tune a compact open-source LLM for customer service summarization to replace GPT-3.5 in production, reducing inference costs and mitigating privacy risks associated with proprietary models.
@@ 4 Error-Focused Summary Evaluation Scheme
To improve the production summarizer and support AI governance, we sampled production inputs and summaries and had subject matter experts and the AI governance team manually review them. Together, we developed a structured evaluation scheme to ensure consistent reviews and systematically surface recurring issues, focusing on identifying summary errors mapped to predefined error types.
Errors fall into seven categories: Content, Entities, Data Elements, Customer Type, Unnecessary Information, Inferred Sentiment, and Language, each with finer-grained sub-labels. Content covers missing, incorrect, or fabricated information in general; Entities covers missing/incorrect identifiers (e.g., item, order, or customer details); Data Elements captures other structured inaccuracies (e.g., dates); Customer Type captures buyer/seller misclassification (WebForm only); Unnecessary Information flags irrelevant/redundant details; Inferred Sentiment flags unsupported emotional inferences; and Language covers language identification and translation errors. The full taxonomy is provided in Table 6 (Appendix).
We rate summaries on a 1–5 scale based on error count and severity: 1 = numerous/critical errors; 2 = one or two errors; 3 = minor error without affecting understanding; 4 = acceptable with no distinct errors but potentially not maximally concise; 5 = error-free.
We use this scheme to evaluate summaries from different LLMs and fine-tuned models (Section 6) and to systematically aanlyze and quantify common GPT-3.5 error patterns (Section 5.2).
@@ 5.1 Data
We collected 10,000 training, 200 development, and 200 test samples per communication channel. All personally identifiable information (PII) in customer interaction log were anonymized by substituting them with artificial fake values.
@@ 5.2 Systematic error analysis
A team of subject matter experts audited GPT-3.5 production summaries (100 BotChat, 68 WebForm, 133 Email, 100 TeammateChat) and labeled errors using our predefined taxonomy (Table 6). We then identified the most frequent error types per channel, summarized in Table LABEL:table:major_error_types.
Table 1: Major error types across communication channels (percentage of total errors per channel).
Error Type BotChat WebForm Email TeammateChat
Unnecessary Information
unn_content_requests_agent 38.0%
unn_content_redundant 29.2% 23.7% 9.9%
unn_content_courtesy 13.0% 33.0%
unn_content_transfer 15.8%
unn_content_webform_email_copy 8.3%
unn_other 7.4%
Content Issues
content_missing 16.5%
content_order 15.5%
content_inaccurate 6.3% 6.3% 5.4%
nothing_to_summarize 7.6%
Entity / Customer Type
entity_missing_item_number 6.3% 8.7%
customer_type_inaccurate 14.6%
Inferred Sentiment
sentiment_inferred_frustrated 24.0%
sentiment_inferred_not_frustrated 6.3%
@@ 5.3 Automatic error-correction
The target errors for auto-correction were selected based on three criteria: (1) frequency, as determined by the systematic error analysis (Section 5.2); (2) manageability, defined as the effectiveness of correcting a given error type using a simple prompting strategy in preliminary experiments; and (3) business criticality, referring to errors whose correction was deemed essential to improve the overall usability and reliability of the generated summaries.
For the BotChat channel, we targeted the error unn_content_requests_agent, defined as unnecessary mentions in summaries that customers requested human assistance, and sentiment_inferred_frustrated and sentiment_inferred_not_frustrated, which capture cases where customer frustration (or lack thereof) was inferred despite not being explicitly stated. For WebForm summaries, we selected unn_content_webform_email_copy, referring to the unnecessary inclusion of customer requests for email copies of submitted webforms. In the TeammateChat channel, we focused on unn_content_transfer and unn_content_courtesy, denoting the unnecessary inclusion of transfer information and courteous exchanges between the teammate and the customer, respectively. For the Email channel, we evaluated unn_content_courtesy in addition to content_order, indicating misleading or incorrect ordering of information in the summary, and unn_content_email_footer, which reflects unnecessary references to email footer content. Finally, all the channels, the error unn_content_redundant, denoting unnecessary redundant information, was targeted for correction under the assumption that increased conciseness would benefit summaries across all channels.
We processed all 40,000 training samples across the four channels using a cascading sequence of error-correction prompts. Each prompt was designed to correct a specific error type, and the prompts were applied sequentially, with the unn_content_redundant correction applied as the final step for all channels. We denote the dataset prior to applying the unn_content_redundant correction as revision version 1 (r1), and the dataset after this correction as revision version 2 (r2). The correction prompts corresponding to each error type are shown in Figure 3 (Appendix).
Notably, the prompt targeting sentiment_inferred_frustrated was intentionally designed to remove all sentiment-related content rather than only inferred frustration. This broader removal does not adversely affect summary quality, as sentiment information is considered non-essential under our evaluation criteria (Section 4).
We employed Llama 3.1 70B as the editor LLM for performing these error corrections. Being an open-source model, Llama 3.1 70B provides reduced privacy risks.
@@ 5.4 LLM fine-tuning
We evaluated a suite of open-source LLMs of varying sizes as student models. Each model was fine-tuned independently using three dataset variants: the original training dataset (org), revision 1 (r1), and revision 2 (r2). Instruction-tuning samples were constructed using the same summarization prompts as our production GPT-3.5 system. Fine-tuning was performed with Low-Rank Adaptation (LoRA) [6], using $\alpha=16$, $r=8$, and dropout =0.1=0.1, applied to all linear modules. To prevent the model from learning to continue the input transcript, we applied prompt loss masking during SFT, computing the language-model loss only over assistant completion tokens (i.e., system and user tokens were excluded from the loss).
[…]
@@ 6 Evaluation
Specifically, we used GPT-4.1 mini and GPT4o as the judge LLMs and performed prompt tuning separately for each communication channel to output ratings on a 1-to-5 scale, based on the original content and corresponding generated summaries. Our automatic evaluation system demonstrated strong correlation with human ratings when benchmarked against other advanced LLM-as-judge methodologies [8, 3]. The Spearman’s $\rho$ [17] and endall’s $\tau$ [11] scores for each channel is shown in Table 2.
Our primary evaluation metric is the mean score of automatic ratings provided by the auto-evaluators on the test dataset.
Table 2: LLM-as-judge alignment to human judgement
Channel $\rho$ $\tau$ Judge LLM
TeammateChat 0.5313 0.4802 GPT4.1 mini
BotChat 0.4735 0.4225 GPT4.1 mini
WebForm 0.4612 0.4203 GPT4.1 mini
Email 0.4551 0.4405 GPT4o